# Fast Quiet-STaR: Thinking Without Thought Tokens

### v1

@@ Abstract
Large Language Models (LLMs) have achieved impressive performance across a range of natural language processing tasks. However, recent advances demonstrate that further gains—particularly in complex reasoning tasks—require more than merely scaling up model sizes or training data. One promising direction is to enable models to “think” during the reasoning process. Recently, Quiet-STaR significantly improves reasoning by generating token-level thought traces, but incurs substantial inference overhead. In this work, we propose Fast Quiet-STaR, a more efficient reasoning framework that preserves the benefits of token-level reasoning while reducing computational cost. Our method introduces a curriculum-learning-based training strategy that gradually reduces the number of thought tokens, enabling the model to internalize more abstract and concise reasoning processes. We further extend this approach to the standard Next Token Prediction (NTP) setting through reinforcement learning-based fine-tuning, resulting in Fast Quiet-STaR NTP, which eliminates the need for explicit thought token generation during inference. Experiments on four benchmark datasets with Mistral 7B and Qwen2.5 7B demonstrate that Fast Quiet-STaR consistently outperforms Quiet-STaR in terms of average accuracy under the same inference time budget. Notably, Fast Quiet-STaR NTP achieves an average accuracy improvement of 9% on Mistral 7B and 5.7% on Qwen2.5 7B, while maintaining the same inference latency. Our code will be available at https://github.com/huangwei200012/Fast-Quiet-STaR.
@@
Wei Huang1 Yizhe Xiong2,3 Xin Ye4 Zhijie Deng5 Hui Chen2,3 Zijia Lin2 Guiguang Ding2,3 1School of Computer Science, Beijing University of Posts and Telecommunications 2Tsinghua University 3Beijing National Research Center for Information Science and Technology (BNRist) 4Kuaishou Technology 5Shanghai Jiao Tong University, Shanghai, China
@@ 2.2 Curriculum Learning
Curriculum Learning is a training strategy that organizes the learning process by first presenting simpler examples and gradually introducing more complex ones. In recent years, curriculum learning has been widely adopted in the training of Large Language Models (LLMs) Xu et al. (2020); Naïr et al. (2024). LDCAL Li et al. (2024) leverages LLMs themselves to assess the difficulty of training instances, guiding the model to learn in an easy-to-hard sequence. TAPIR Yue et al. (2024) constructs a task-aware curriculum scheduling framework that dynamically adjusts the task distribution and progressively increases task complexity. Moreover, curriculum learning has also been employed to improve LLMs’ understanding of long contexts by gradually increasing the context window size during training Grattafiori et al. (2024b). Kimi 1.5 Team et al. (2025) integrates a curriculum learning strategy during the reinforcement learning stage, allowing the model to start with simpler question before transitioning to more complex ones.
@@ 3 Methodology
The training procedure of Fast Quiet-STaR Zelikman et al. (2024) is illustrated in Figure 2. Building upon Quiet-STaR, we propose a progressive, multi-stage training framework inspired by the principles of curriculum learning. This approach facilitates a gradual transition from easy to hard reasoning paradigms. In particular, during the final stage of training, we incorporate reinforcement learning to transition the reasoning paradigm of Quiet-STaR to the the standard NTP paradigm.
@@ 4.1 Experimental Settings
Implementation details. All training experiments are conducted on 8 H800 GPUs. For Quiet-STaR, we train for 100 steps, and for Fast Quiet-STaR, we select the last ckpt in the previous stage for initialization and train for another 50 steps for each training stage. See Appendix A for more details.
@@ 4.3.1 Ablation Studies
Figure 3: Comparison of the average accuracy of Fast Quiet-STaR and UP Quiet-STaR with training steps.
@@
Fast Quiet-STaR NTP Without Curriculum Learning. To evaluate the effectiveness of the curriculum learning procedure “16-8 → 12-4 → 8-4 → NTP”, we omit the intermediate stages. Specifically, we directly initialize the pre-trained model with Quiet-STaR 16-8 and use its log-likelihood loss as the reference for computing rewards in reinforcement learning, resulting in Quiet-STaR NTP. As shown in Table 3, this shortcut results in a 5.2% drop in average accuracy compared to Fast Quiet-STaR NTP obtained through the full curriculum. These results underscore the critical role of the curriculum learning process in enhancing overall model performance.
@@ 4.3.4 Thought token analysis
We visualize the thought tokens generated by Quiet-STaR 8-4 and Fast Quiet-STaR 8-4 at key positions—tokens most informative for final predictions—on the GSM8K dataset to examine their internal reasoning behavior. As shown in Figure 6, Quiet-STaR 8-4 produces relatively unstructured thoughts, indicating incomplete acquisition of the Quiet-STaR reasoning paradigm. In contrast, Fast Quiet-STaR 8-4 demonstrates more abstract and goal-directed reasoning behavior. These observations indicate that the incorporation of a curriculum learning strategy—progressing from easier to more difficult—enables Fast Quiet-STaR to gradually acquire the ability to perform effective reasoning under resource constraints.
@@ Appendix A Implementation details
Our experimental settings are consistent with Quiet-STaR Zelikman et al. (2024). Specifically, we employ the AdamW optimizer with a warm-up step count of 20, a weight decay of 0.001, and a batch size of 8. For Quiet-STaR, we train for 100 steps, and for Fast Quiet-STaR, we select the last ckpt in the previous stage for initialization and train for 50 steps for each stage. The learning rates are adjusted slightly depending on the model: we use a learning rate of 1e-6 for Mistral Jiang et al. (2023) and 8e-6 for Qwen2.5 Qwen et al. (2025). During training, we perform sampling with a temperature of T=1. For evaluation, we adopt greedy decoding to ensure deterministic outputs. All training experiments are conducted on eight H800 GPUs. For measuring the Time to First Token (TTFT), we utilize a single H800 GPU and fix the context length to 256. TTFT is defined as the elapsed time between the moment the model receives the full input sequence and the generation of the first token.

### v2

@@ Abstract
Large Language Models (LLMs) have achieved impressive performance across a range of natural language processing tasks. However, recent advances demonstrate that further gains—particularly in complex reasoning tasks—require more than merely scaling up model sizes or training data. One promising direction is to enable models to “think” during the reasoning process. Recently, Quiet-STaR significantly improves reasoning by generating token-level thought traces, but incurs substantial inference overhead. In this work, we propose Fast Quiet-STaR, a more efficient reasoning framework that preserves the benefits of token-level reasoning while reducing computational cost. Our method introduces a curriculum-learning-based training strategy that gradually reduces the number of thought tokens, enabling the model to internalize more abstract and concise reasoning processes. We further extend this approach to the standard Next Token Prediction (NTP) setting through reinforcement learning-based fine-tuning, resulting in Fast Quiet-STaR NTP, which eliminates the need for explicit thought token generation during inference. Experiments on four benchmark datasets with Mistral 7B and Qwen2.5 7B demonstrate that Fast Quiet-STaR consistently outperforms Quiet-STaR in terms of average accuracy under the same inference time budget. Notably, Fast Quiet-STaR NTP achieves an average accuracy improvement of 9% on Mistral 7B and 5.7% on Qwen2.5 7B, while maintaining the same inference latency.
@@
Wei Huang1thanks: These authors contributed equally to this work., Yizhe Xiong2,311footnotemark: 1, Xin Ye411footnotemark: 1, Zhijie Deng5, Hui Chen2,3, Zijia Lin2, Guiguang Ding2,3 1School of Computer Science, Beijing University of Posts and Telecommunications, 2Tsinghua University, 3Beijing National Research Center for Information Science and Technology (BNRist), 4Kuaishou Technology, 5Shanghai Jiao Tong University, Shanghai, China Correspondence: jichenhui2012@gamil.com, linzijia07@tsinghua.org.cn
@@ 2.2 Curriculum Learning
Curriculum Learning is a training strategy that organizes the learning process by first presenting simpler examples and gradually introducing more complex ones. In recent years, curriculum learning has been widely adopted in the training of Large Language Models (LLMs) Xu et al. (2020); Na¨ır et al. (2024). LDCAL Li et al. (2024) leverages LLMs themselves to assess the difficulty of training instances, guiding the model to learn in an easy-to-hard sequence. TAPIR Yue et al. (2024) constructs a task-aware curriculum scheduling framework that dynamically adjusts the task distribution and progressively increases task complexity. Moreover, curriculum learning has also been employed to improve LLMs’ understanding of long contexts by gradually increasing the context window size during training Grattafiori et al. (2024b). Kimi 1.5 Team et al. (2025) integrates a curriculum learning strategy during the reinforcement learning stage, allowing the model to start with simpler question before transitioning to more complex ones.
@@ 3 Methodology
The training procedure of Fast Quiet-STaR is illustrated in Figure 2. Building upon Quiet-STaR, we propose a progressive, multi-stage training framework inspired by the principles of curriculum learning. This approach facilitates a gradual transition from easy to hard reasoning paradigms. In particular, during the final stage of training, we incorporate reinforcement learning to transition the reasoning paradigm of Quiet-STaR to the the standard NTP paradigm.
@@ 4.1 Experimental Settings
Implementation details. All training experiments are conducted on 8 H800 GPUs. For Quiet-STaR, we train for 100 steps, and for Fast Quiet-STaR, we select the last checkpoint in the previous stage for initialization and train for another 50 steps for each training stage. See Appendix A for more details.
@@ 4.3.1 Ablation Studies
Figure 3: Comparison of the average accuracy of Fast Quiet-STaR and Rev Quiet-STaR with training steps.
@@
Fast Quiet-STaR NTP Without Curriculum Learning. To evaluate the effectiveness of the curriculum learning procedure “16-8 → 12-4 → 8-4 → NTP”, we omit the intermediate stages. Specifically, we directly initialize the pre-trained model with Quiet-STaR 16-8 and use its log-likelihood loss as the reference for computing rewards in reinforcement learning, resulting in Quiet-STaR NTP. As shown in Table 3, this shortcut results in a 5.2% drop in average accuracy compared to Fast Quiet-STaR NTP obtained through the full curriculum. These results underscore the critical role of the curriculum learning process in enhancing the overall performance of the model.
@@ 4.3.3 Performance on Generation Tasks
4.3.4 Analysis of long-term inference savings
The training process of Fast Quiet-STaR is highly efficient. The entire training pipeline (from Quiet-STaR 16-8 → Fast Quiet-STaR 12-4 → Fast Quiet-STaR 8-4 → Fast Quiet-STaR NTP) requires only 0.5M tokens for continue pre-training. This process takes just 54 minutes on 8 H800 GPUs. We use prompts containing 256 tokens and allow the model to generate 128 tokens for demonstration and analysis purposes. On a single H800 GPU, the inference latency of Fast Quiet-STaR-NTP is 3.2 seconds, compared to 52.7 seconds for Quiet-STaR 16-8(more scenarios with inference latency are in Table tab:time). For 67 end-to-end inference runs, Quiet-STaR 16-8 requires approximately 59 minutes in total, whereas Fast Quiet-STaR-NTP completes the same task in just 4 minutes—yielding an overall speedup of 55 minutes. In other words, the time saved from just 67 inference runs is sufficient to offset the entire training cost of Fast Quiet-STaR-NTP. More importantly, in real-world deployments, models are typically required to perform millions of inference runs—far exceeding 67—where our approach would demonstrate substantial advantages in total inference cost.
@@ 4.3.5 Thought token analysis
We visualize the thought tokens generated by Quiet-STaR 8-4 and Fast Quiet-STaR 8-4 at key positions—tokens most informative for final predictions—on the GSM8K dataset to examine their internal reasoning behavior. As shown in Figure 6, Quiet-STaR 8-4 produces relatively unstructured thoughts, indicating incomplete acquisition of the Quiet-STaR reasoning paradigm. In contrast, Fast Quiet-STaR 8-4 demonstrates more abstract and goal-directed reasoning behavior. These observations indicate that the incorporation of a curriculum learning strategy—progressing from easier to more difficult—enables Fast Quiet-STaR to gradually acquire the ability to perform effective reasoning under resource constraints. For more examples, please refer to Appendix B.
@@ 5 Conclusion
Limitations
Despite the notable improvements in inference efficiency and performance achieved by Fast Quiet-STaR, several limitations remain. First, the evaluation in this study primarily focuses on mathematical and logical reasoning tasks, leaving its generalization capability to other domains yet to be thoroughly validated. Second, this method is only for the Quiet-STaR reasoning method.
@@ Appendix A Implementation details
Our experimental settings are consistent with Quiet-STaR Zelikman et al. (2024). Specifically, we employ the AdamW optimizer with a warm-up step count of 20, a weight decay of 0.001, and a batch size of 8. For Quiet-STaR, we train for 100 steps, and for Fast Quiet-STaR, we select the last checkpoint in the previous stage for initialization and train for 50 steps for each stage. The learning rates are adjusted slightly depending on the model: we use a learning rate of 1e-6 for Mistral Jiang et al. (2023) and 8e-6 for Qwen2.5 Qwen et al. (2025). During training, we perform sampling with a temperature of T=1. For evaluation, we adopt greedy decoding to ensure deterministic outputs. All training experiments are conducted on eight H800 GPUs. For measuring the Time to First Token (TTFT), we utilize a single H800 GPU and fix the context length to 256. TTFT is defined as the elapsed time between the moment the model receives the full input sequence and the generation of the first token.
We report the version numbers of used packages in Table 5.
Figure 7: More examples of the text and its thought process at key positions.
Package Version Package Version
PyTorch 2.1.0 transformers 4.46.0
deepspeed 0.10.0 tokenizers 0.13.3
datasets 2.14.3
Table 5: Versions of used packages.
Appendix B More Thought token analysis
Based on the settings described in Section 4.3.5, we further visualize examples of thought tokens generated at key positions by Quiet-STaR 8-4 and Fast Quiet-STaR 8-4, as illustrated in Figure 7. Compared to Quiet-STaR 8-4, the thought tokens generated by Fast Quiet-STaR 8-4 exhibit a more abstract and directive reasoning process.
Appendix C License for Scientific Artifacts
The Open web math Paster et al. (2023) is licensed under ODC-By 1.0 License [^1] https://spdx.org/licenses/ODC-By-1.0.html. The Mistral model Jiang et al. (2023) and Qwen2.5 model Qwen et al. (2025) is licensed under Apache License 2.0 license [^2] https://choosealicense.com/licenses/apache-2.0/. The evaluation datasets Cobbe et al. (2021); Talmor et al. (2018); Sap et al. (2019); Bisk et al. (2020) are subject to the MIT license [^3] https://choosealicense.com/licenses/mit/. All usages of scientific artifacts in this paper obey the corresponding licenses.